OpenStudy (sambam):

Please Solve 4^x = 256 also if you can please show your work so I can actually learn how to do it.

OpenStudy (anonymous):

Do you know about logarithms?

OpenStudy (sambam):

Not really

OpenStudy (anonymous):

\[ 4^x = 4^4\\ x=4\]

OpenStudy (anonymous):

That's just one way. As it turns out, 256 is a power of 4.

OpenStudy (anonymous):

no need for logs

OpenStudy (anonymous):

LIke he said.

OpenStudy (anonymous):

I know she does not need logs, but if she knew about them that would be good. Unless you are highly familiar with all the powers of a number, it's just a calculator game until you hit the right power.

OpenStudy (majacw):

WolframAlpha.com and Mathway are great (FREE) math sites. Good luck!

OpenStudy (skullpatrol):

What is on the left side of this equation: $$4^x = 256 $$and what is on the right side of the equation look very different correct?

OpenStudy (luigi0210):

I prefer Luis's method over logs

OpenStudy (skullpatrol):

@sambam Try to make the sides "look" as much as possible the same.

OpenStudy (jhannybean):

\[\large 4^x = 256 \]\[\large 256= 4^4\]\[\large 4^x = 4^4\]\[\large \cancel4^x = \cancel4^4\]\[\large x= 4\] when you have teh same base, you can just eliminate them and evaluate the powers.

OpenStudy (nincompoop):

4^x = 256 Perform the prime factorization of the left hand side. 4^x = 2^(2 x): 2^(2 x) = 256 Perform the prime factorization of the right hand side. 256 = 2^8: 2^(2 x) = 2^8 Equate exponents. Equate exponents of 2 on both sides: 2 x = 8 Solve the linear equation. All equations give x = 4 as the solution: Answer: | x = 4

OpenStudy (anonymous):

I agree Luigi, but if you are not aware of the powers of 4 in this case you could just be sitting there. These ease of solution comes from this knowledge. I thought the questioner might not be aware of the connection between 4 and 256.

OpenStudy (anonymous):

I get on board with nincompoop. There's some math there and not so much luck.

OpenStudy (nincompoop):

This is why it is important to have a very strong grasp of the fundamentals. One can always exploit solutions from the basics.

OpenStudy (skullpatrol):

Or just divide 256 by 4 :)

OpenStudy (anonymous):

Preach on nincompoop! Maybe you have been misnamed?

OpenStudy (anonymous):

I believe sambam left the arena. Just a bunch of math - o - philes here now.